OpenStudy (anonymous):

f(x+1)^1 f(x)= 3x+2 I did it, but I have a question about the square root??

OpenStudy (anonymous):

sorry, not square root, square

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

i did it like this f(x+1)^1 = 3(x^2+2x+1) + 2, but my teacher did it like this [3(x+1)+2]^2, why did he do that?

OpenStudy (anonymous):

what are you trying to find in that equation? why is f(x+1) raised to the power of 1?

OpenStudy (anonymous):

i did it like this f(x+1)^1 = 3(x^2+2x+1) + 2, but my teacher did it like this [3(x+1)+2]^2, why did he do that?

OpenStudy (anonymous):

sorry, i meant power of two

OpenStudy (anonymous):

so your question is what is f(x+1)^2 when f(x) = 3x+2?

OpenStudy (anonymous):

no, my question is why did my teacher put the ^2 on the outside of the equation, like this [3(x+1)+2]^2

OpenStudy (anonymous):

if that is what you mean than what you do is you substitute x with x+1 and then you square it, so f(x+1) = 3(x+1)+2, and since you have f(x+1)^2 it means you are squaring the entire function so you have (3(x+1)+2)^2

OpenStudy (anonymous):

oh ok. I thought I was only squaring (x+1)

OpenStudy (anonymous):

f(x) is one term so f(x)^2 = f(x)*f(x)

OpenStudy (anonymous):

oh ok. thx